# A Unified Pulse-Shaped OFDM Framework for Chirp-Domain Waveforms: Continuous-Time Modeling and Practical I/O Analysis

### 15 Mar 2026 v1

@@ V-B Practical Waveform based I/O Relation
The passband counterpart of (41) is given by
$y_{pb}(t)=$ $\sum_{p=1}^{P}\tilde{h}_{p}\cos(2\pi(f_{c}+\nu_{p})(t-\tau_{p}))x(t-\tau_{p})$
$+y_{pb}^{I}(t)+z_{pb}(t),$ (48)
where $\tilde{h}_{p}$ is a real-valued gain and $h_{p}=\tilde{h}_{p}e^{-j2\pi f_{c}\tau_{p}}$, $y_{pb}^{I}(t)$ denotes other bandpass signals or interference, and $z_{pb}(t)$ is the bandpass noise. To obtain $y(t)$, $y_{pb}(t)$ must first be down-converted and then passed through a low-pass filter to reject out-of-band noise and signals including $y_{pb}^{I}(t)$ and the high-frequency component around $2f_{c}$. According to the established communication theory[13, 14, 15], the noise $z(t)$ is modeled as having infinite bandwidth and therefore infinite power. In other words, $y(t)$ in (41) is an idealized model that has not yet accounted for the effects of low-pass or receive filtering.
Substituting $x_{\rm AF}^{(a)}(t)$ in (36) as $x(t)$ into (41) and comparing with (42), one can see that the time-varying channel tap in (43) is obtained by directly sampling $y(t)$ with an interval of $\frac{T}{N}$, where the pulse-shaping filter $a(t)$ is a unit-amplitude Nyquist pulse. Therefore, the effect of practically required low-pass or receive filtering is not considered in (42), and the statistical characteristics of of $\mathbf{z}$ may not be well-justified.
Recall that the approximate aliased chirps in (38) are orthogonal to each other when $a(t)$ is a root-Nyquist pulse-shaping filter. Maintaining this orthogonality at the output of the receive filter under an ideal channel naturally suggests employing the same root Nyquist pulse-shaping filter as the receive filter, which also serves as the matched filter.
Let us apply the matched filter $a^{*}(-t)$ to $y(t)$. The sampled output of the matched filter is given by
@@
$h_{k^{\prime},l}^{\rm MF}=\sum_{p=1}^{P}h_{p}\,e^{j2\pi\nu_{p}(k^{\prime}-l)\frac{T}{N}}\,\mathcal{A}_{a}(l\frac{T}{N}+\tau_{1}-\tau_{p},\nu_{p}),$ (50)
@@
We can now see that, in a practical receiver, the time-varying channel tap in (50) does not have the elegant form assumed in (43), even with on-grid delays, as $\mathcal{A}_{a}(l\frac{T}{N},\nu_{p})$ is generally not zero for $\nu_{p}\neq 0$. In other words, the effective channel cannot be represented by a pure DD spreading function. As a result, the ideal I/O relation in (47) does not exist in practice, thereby rendering the subsequent discussion and conclusions based on it less relevant to practical systems.
Similarly, after discarding the CPP, (49) can be written in matrix form as
@@
where $\mathbf{z}^{\rm MF}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$ since $a(t)$ is a root-Nyquist pulse for the Nyquist interval $\frac{T}{N}$ with unit energy $\int|a(t)|^{2}dt=1$. Further applying the DAFT matrix $\mathbf{A}$ to $\mathbf{y}^{\rm MF}$, we have
@@
(52) shows the explicit practical I/O relation in the DAFT domain. In particular, $\mathbf{H}_{u}^{\rm MF}$ represents the effective coupling coefficient from the transmitted DAFT-domain symbol $X[n]$ to the observed output symbol $Y[n^{\prime}]$ after pulse shaping, physical channel propagation, matched filtering, sampling, and the folding of the CPP portion. This relation provides a foundation for pilot design and channel equalization for practical AFDM waveforms.
@@ VI Simulation Results
In the simulations, the AFDM parameters are set to $T=266.667\mu s$, $N=1024$, and the corresponding subcarrier spacing is $\frac{1}{T}=3.75$ kHz. The transmitted symbols are drawn from a $16$-QAM constellation with average power $\sigma_{X}^{2}=1$. To generate the approximate aliased AFDM waveform, the AFDM sequence is pulse-shaped by an square root raised cosine (SRRC) filter with roll-off factor $\beta$ and span $Q\frac{T}{N}$. The multipath channel follows the standard EVA profile [23], consisting of $P=9$ paths with delays that are off-grid with respect to $\frac{T}{N}$. The carrier frequency is $f_{c}=5$GHz, and the Doppler frequency of the $p$-th path is generated using Jakes’ formula $\nu_{p}=\nu_{\textrm{max}}\cos(\theta_{p})$, where the maximum Doppler $\nu_{\textrm{max}}$ is determined by the terminal speed and $\theta_{p}$ is uniformly distributed over $\pi,\pi]$.
@@ VI-B Verification of I/O relation
Next, we verify the derived I/O relation for implemented AFDM waveforms over DD channels, where the chirp parameters are $c_{1}=1$, $c_{2}=2$, respectively. The accuracy of the theoretical model is evaluated in terms of the normalized mean square error (NMSE) between the received DAFT domain signal $\mathbf{y}^{\rm MF}_{u}$ predicted by the derived I/O relation and the received signal obtained from direct waveform simulation through the continuous-time DD channel.
@@ VI-B1 Impact of Terminal Speed
Fig. 2 shows the NMSE versus terminal speed, which determines the maximum Doppler shift in the DD channel. The terminal speed is varied from 0 to $500$ km/h, while the roll-off factor and filter span are fixed at $\beta=0.2$ and $Q=12$, respectively. The results demonstrate that the derived I/O relation accurately predicts the received signal over a wide range of Doppler conditions. In particular, the NMSE remains consistently below 35-35dB even at high mobility. This indicates that variations in terminal speed do not significantly affect the consistency between the theoretical model and the waveform simulation, confirming that the derived I/O relation maintains a high level of accuracy under highly time-varying channel conditions.
@@ VI-B2 Impact of Roll-Off Factor
Fig. 3 presents the NMSE performance under different roll-off factors of the sample-wise pulse-shaping filter used in the implemented waveform. The roll-off factor is varied from $0.1$ to $0.9$, while the terminal speed and filter span are fixed at $500$ km/h and $Q=12$, respectively. The results show that the derived I/O relation remains valid across various roll-off factors. In particular, the NMSE consistently remains within the relatively low range of 37-37 dB to 33-33 dB as the roll-off factor varies. Although the roll-off factor has some influence on the error level, its overall impact is limited. These results indicate that the derived I/O relation can still accurately characterize the I/O behavior of the implemented AFDM waveform under different pulse-shaping settings.
@@ VI-B3 Impact of Filter Span
Fig. 4 illustrates the NMSE as a function of the filter span, where $Q$ is varied from $6$ to $20$, while the terminal speed and roll-off factor are fixed at $500$ km/h and $\beta=0.2$, respectively. The results show that the NMSE remains within the range of 36-36 dB to 32-32 dB while exhibiting a clear decreasing trend as the filter span increases. This indicates that a larger filter span provides a more accurate approximation to the ideal pulse-shaping filter $a(t)$, thereby reducing the approximation error and improving the agreement between the theoretical model and the direct waveform simulation results. Moreover, when the span becomes sufficiently large, the improvement in NMSE gradually diminishes, suggesting that the error reduction begins to saturate. These results further confirm that the derived I/O relation accurately captures the behavior of implemented AFDM waveforms.
@@ VI-C Implementation Complexity Comparison
We compare the implementation complexity of AFDM with that of ODDM. Note that the implemented AFDM signal has a time span of approximately $T$ and a bandwidth of approximately $\frac{N}{T}$. Then, the parameters of ODDM are configured as follows:
  • Number of subcarriers: $N_{\rm OD}$
  • Number of ODDM symbols: $M_{\rm OD}=\frac{N}{N_{\rm OD}}$
  • Subcarrier spacing: $\frac{1}{T}$
to occupy nearly the same TF region.
@@
As shown above, AFDM is an $N$-subcarrier system, and the complexity associated with the inverse fast Fourier transform (IFFT) operation is $\mathcal{O}(N\log N)$. In contrast, ODDM is an $N_{\rm OD}$-subcarrier system and requires $M_{\rm OD}$ IFFTs of size $N_{\rm OD}$, resulting in a complexity of
$M_{\rm OD}\mathcal{O}(N_{\rm OD}\log N_{\rm OD})=\mathcal{O}(N\log N_{\rm OD})\ll\mathcal{O}(N\log N),$
because usually $N_{\rm OD}\ll N$. Furthermore, it is known that a CP is not necessary for ODDM signals [12], as ODDM is essentially a pseudo-impulse-based transmission scheme for DD channels [9]. As a result, the implementation complexity of ODDM is significantly lower than that of AFDM.
@@ VII Conclusion
This paper develops a unified framework for chirp-domain waveforms, including OCDM and AFDM. By examining their continuous-time representations, it has been shown that these waveforms naturally fall within the conventional WH framework for multicarrier waveforms, where the root chirp corresponds to the prototype pulse. This perspective further reveals that chirp-domain waveforms can be interpreted as PS-OFDM with a constant-envelope prototype pulse, under which their PSD can be characterized analytically.
Practical implementations based on sub-Nyquist time-domain samples have also been analyzed. The resulting aliased waveforms and the orthogonality properties of the embedded aliased chirps have been investigated, showing that orthogonality can still be preserved in implemented waveforms through appropriate sample-wise pulse shaping. An exact I/O relation over DD channels has been derived, revealing that the effective channel observed at practical receivers does not generally follow the conventional DD spreading-function model. Simulations are provided to verify the derived theoretical results. Future work may extend the proposed framework to receiver design and waveform optimization under off-grid DD channels, as well as to multi-antenna systems.
@@ Appendix A Orthogonality Analysis of Aliased Chirps
$\hat{\phi}_{n}(t)=e^{j2\pi\left(c_{2}n^{2}+\frac{c_{1}}{\Delta t^{2}}t^{2}+\frac{n}{T}t-\frac{q_{n}(t)}{\Delta t}t\right)}$ (53)
where $T=N\Delta t$, $C=2N|c_{1}|$, and the interval index $q_{n}(t)$ updates at boundaries $t_{n,q}=\frac{T}{C}q-\frac{n}{C}\Delta t$. This boundary condition allows us to express $q_{n}(t)$ explicitly using the floor function:
@@
We then partition the integration interval $[0,T]$ into $C$ segments of length $T/C$ and introduce a local time variable $\varepsilon\in[0,1)$ and an interval index $c\in\{0,1,\dots,C-1\}$, such that $t=(\varepsilon+c)\frac{T}{C}$.
Substitute this into $\Delta q(t)$. Because $c$ is an integer, we can use the floor function property $\lfloor x+c\rfloor=\lfloor x\rfloor+c$ to factor $c$ out:
$\Delta q(t)$ $=\left\lfloor(\varepsilon+c)+\frac{n}{N}\right\rfloor-\left\lfloor(\varepsilon+c)+\frac{n^{\prime}}{N}\right\rfloor$
$=\left(\left\lfloor\varepsilon+\frac{n}{N}\right\rfloor+c\right)-\left(\left\lfloor\varepsilon+\frac{n^{\prime}}{N}\right\rfloor+c\right)$
@@
The integer $c$ completely cancels out of the expression. Because the resulting expression depends solely on the local time variable $\varepsilon$ and is independent of $c$, we introduce the shorthand $\equiv\Delta q(\varepsilon)$. This explicitly denotes that $\Delta q$ is periodic across every segment $c$.
Substituting $t=(\varepsilon+c)\frac{T}{C}$ into the integral and rewriting as a sum over $c$, we obtain
$I_{n,n^{\prime}}\propto\frac{T}{C}\sum_{c=0}^{C-1}\int_{0}^{1}e^{j2\pi\left(\frac{n^{\prime\prime}}{T}-\frac{\Delta q(\varepsilon)}{\Delta t}\right)(\varepsilon+c)\frac{T}{C}}d\varepsilon.$ (58)
Using $T=N\Delta t$, the phase simplifies to $\frac{n^{\prime\prime}-N\Delta q(\varepsilon)}{C}(\varepsilon+c)$. We define the integer function $\eta(\varepsilon)=n^{\prime\prime}-N\Delta q(\varepsilon)$, allowing us to separate the variables:
$I_{n,n^{\prime}}\propto\frac{T}{C}\int_{0}^{1}e^{j2\pi\frac{\eta(\varepsilon)}{C}\varepsilon}\underbrace{\left[\sum_{c=0}^{C-1}e^{j2\pi\frac{\eta(\varepsilon)}{C}c}\right]}_{\text{Geometric Sum }S(\varepsilon)}d\varepsilon,$ (59)

### 31 Mar 2026 v2

@@ V-B Practical Waveform based I/O Relation
In practice, the received real-valued passband signal is
$y_{pb}(t)=y_{pb}^{S}(t)+y_{pb}^{I}(t)+z_{pb}(t),$ (48)
where the desired signal component is
$y_{pb}^{S}(t)$ $=\sqrt{2}\sum_{p=1}^{P}\tilde{h}_{p}\cos\!\left(2\pi(f_{c}+\nu_{p})(t-\tau_{p})\right)x(t-\tau_{p})$
$=\Re\left\{y^{S}(t)\sqrt{2}e^{j2\pi f_{c}t}\right\},$ (49)
and $y^{S}(t)=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(t-\tau_{p})}x(t-\tau_{p})$ is its complex baseband equivalent with respect to the system carrier reference $\sqrt{2}e^{j2\pi f_{c}t}$. Here, $\tilde{h}_{p}$ is a real-valued path gain, $h_{p}=\tilde{h}_{p}e^{-j2\pi f_{c}\tau_{p}}$ is the corresponding complex baseband gain, $y_{pb}^{I}(t)$ denotes the out-of-band interference component, and $z_{pb}(t)$ is passband AWGN with a two-sided PSD of $\frac{N_{0}}{2}$.
To obtain the baseband signal, $y_{pb}(t)$ is first filtered by a real-valued bandpass filter (BPF) $r_{pb}(t)$ to remove $y^{I}_{pb}(t)$ and is then down-converted[26]. Let $r(t)$ denote the baseband equivalent of $r_{pb}(t)$. After mixing, we obtain
$\tilde{y}(t)=$ $y^{S}(t)\star r(t)+(y^{S}(t)\star r(t))^{*}e^{-j4\pi f_{c}t}$
$+z(t)\star r(t)+(z(t)\star r(t))^{*}e^{-j4\pi f_{c}t},$ (50)
where $z(t)$ is the baseband equivalent of $z_{pb}(t)$. Further applying a low-pass filter (LPF) $\tilde{r}(t)$ to remove the components around $2f_{c}$ yields
$y(t)$ $=y^{S}(t)\star r(t)\star\tilde{r}(t)+z(t)\star r(t)\star\tilde{r}(t),$ (51)
where the joint effect of the BPF and the LPF, given by $r(t)\star\tilde{r}(t)$, can be interpreted as an equivalent baseband receive filter[9, 27].
In classical communication theory [16, 17, 18], the baseband noise $z(t)$ is often modeled as having infinite bandwidth and, consequently, infinite power for mathematical convenience. To justify this model, a subsequent filtering stage is typically introduced, which aligns exactly with (51). Comparing (41) with (51), it is clear that (41) represents an idealized model that neglects the practical receive filtering mentioned above.
Substituting (36) into (41), it can be seen that the time-varying channel tap $h_{k^{\prime},l}$ in (42) is a superposition of pure path-wise DD components, obtained by sampling $y(t)$ with interval $\frac{T}{N}$ for a unit-amplitude Nyquist pulse $a(t)$. Therefore, the effect of inevitable receive filtering is not captured in (42), and the statistical characterization of $z[k^{\prime}]$ may not be well justified.
Recall that the approximate aliased chirps in (38) are orthogonal to each other when $a(t)$ is a root-Nyquist pulse-shaping filter. Maintaining this orthogonality at the output of the receive filter under an ideal channel naturally suggests employing the same root-Nyquist pulse-shaping filter as the receive filter, which also serves as the matched filter and the anti-aliasing filter.
Let us apply the matched filter $a^{*}(-t)$ to $y(t)$, and let $\mathcal{A}_{a,a}(\cdot)$ denote the ambiguity function of $a(t)$. The sampled output of the matched filter is given by
@@
$\hskip-8.53581pth_{k^{\prime},l}^{\rm MF}=\sum_{p=1}^{P}h_{p}\,e^{j2\pi\nu_{p}(k^{\prime}\frac{T}{N}-\tilde{\tau}_{p})}\,\mathcal{A}_{a,a}(l\frac{T}{N}-\tilde{\tau}_{p},-\nu_{p}),$ (53)
@@
Note that for other pairs of transmit and receive pulses/filters, the equivalent channel taps can be obtained by replacing $\mathcal{A}_{a,a}(\cdot)$ in (53) with the appropriate cross-ambiguity function.
We can now see that, in a practical receiver, the time-varying channel tap in (53) generally does not take the simplified form assumed in (43), even with on-grid delays, because the term $\mathcal{A}_{a,a}(l\frac{T}{N},-\nu_{p})$ is typically non-zero for $\nu_{p}\neq 0$. Consequently, the effective channel generally cannot be expressed as a superposition of pure path-wise DD components. This implies that the I/O relation in (47) generally does not hold in practice, which may limit the relevance of discussions and conclusions based solely on that model.
Remark 3
For LTI channels with $\nu_{p}=0$ for all $p$, the channel model in (43) remains valid without explicitly accounting for the receive filter, as its effect can be fully absorbed into the equivalent on-grid channel. Provided that the equivalent channel remains LTI, which is typically unknown and must be estimated using pilot signals in practice, all analyses and conclusions based on the on-grid LTI channel assumption continue to hold.
Similarly, after discarding the CPP and letting $\mathbf{y}^{\rm MF}=[y^{\rm MF}[0],\cdots,y^{\rm MF}[N-1]]^{\mathsf{T}}$, (52) can be written in matrix form as
@@
where $\mathbf{z}^{\rm MF}=[z^{\rm MF}[0],\cdots,z^{\rm MF}[N-1]]^{\mathsf{T}}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$ because $a(t)$ is a root-Nyquist pulse for the Nyquist interval $\frac{T}{N}$ and has unit energy $\int|a(t)|^{2}dt=1$. Further applying the DAFT matrix $\mathbf{A}$ to $\mathbf{y}^{\rm MF}$, we have
@@
(55) provides the explicit practical I/O relation in the DAFT domain. Specifically, the effective coupling coefficients between $\mathbf{x}_{u}$ and $\mathbf{y}_{u}^{\rm MF}$, represented by $\mathbf{H}_{u}^{\rm MF}$, account for the entire signal chain, including the IDAFT transformation, pulse shaping, physical channel propagation, receive filtering, sampling, CPP folding, and the final DAFT transformation. Since this relation exhibits a non-negligible deviation from $\mathbf{H}_{u}$ in (47), (55) serves as a necessary foundation for pilot design and channel equalization in practical AFDM systems.
@@ VI Simulation Results
[…]
@@ VI-B Verification of I/O relation
Next, we verify the derived I/O relation for implemented AFDM waveforms over DD channels, using the same chirp parameters $c_{1}=\frac{1}{4N}$ and $c_{2}=\frac{1}{3N}$. The accuracy of the theoretical model is evaluated in terms of the normalized mean square error (NMSE) between the received DAFT domain signal $\mathbf{y}^{\rm MF}_{u}$ predicted by the derived I/O relation and the received signal obtained from direct waveform simulation through the continuous-time DD channel.
@@ VI-B1 Impact of Terminal Speed
Fig. 2 shows the NMSE versus terminal speed, which determines the maximum Doppler shift in the DD channel. The terminal speed is varied from 0 to $500$ km/h, while the roll-off factor and filter span are fixed at $\beta=0.2$ and $Q=12$, respectively. The results demonstrate that the derived I/O relation accurately predicts the received signal over a wide range of Doppler conditions. In particular, the NMSE remains consistently below 50-50 dB even at high mobility. This indicates that variations in terminal speed do not significantly affect the consistency between the theoretical model and the waveform simulation, confirming that the derived I/O relation maintains a high level of accuracy under highly time-varying channel conditions.
@@ VI-B2 Impact of Roll-Off Factor
Fig. 3 shows the NMSE performance across various roll-off factors for the pulse-shaping filter used in the implemented waveform. With the terminal speed and filter span fixed at $500$ km/h and $Q=12$, respectively, the NMSE ranges from 39-39 dB to 62-62 dB. The results exhibit a clear decreasing trend as the roll-off factor increases, indicating that, for a fixed filter span, the pulse-shaping filter $a(t)$ with a larger roll-off factor can be approximated more accurately. This reduces truncation error and improves the agreement between the theoretical model and the waveform simulation. Moreover, as the roll-off factor becomes sufficiently large, the improvement in NMSE gradually diminishes, suggesting that the error reduction begins to saturate. These results demonstrate that the derived I/O relation accurately characterizes the behavior of AFDM waveforms under different pulse-shaping settings.
@@ VI-B3 Impact of Filter Span
Fig. 4 illustrates the NMSE as a function of the filter span $Q$, which is varied from $6$ to $20$, with the terminal speed and roll-off factor fixed at $500$ km/h and $\beta=0.2$, respectively. The NMSE decreases from 40-40 dB to 57-57 dB as $Q$ increases, since a larger span provides a more accurate approximation of $a(t)$ for a fixed roll-off factor. Similar to the previous observation, for sufficiently large filter span, the improvement in NMSE becomes marginal, indicating diminishing returns in error reduction. These results further confirm that the derived I/O relation accurately captures the behavior of implemented AFDM waveforms under different truncation lengths.
@@ VI-C Implementation Complexity Comparison
We compare the implementation complexity of AFDM with that of ODDM. Note that the implemented AFDM signal occupies a time-bandwidth product of approximately $T\times\frac{N}{T}=N$. This provides $N$ degrees of freedom (DoF) to carry $N$ digital symbols, which is consistent with classical communication theory [2, 18, 29]. This also explains why AFDM is implemented using sub-Nyquist samples: the ideal waveform occupies excessive bandwidth for transmitting only $N$ symbols, leading to spectral efficiency well below the available DoF. For a fair comparison, the number of subcarriers $N_{\rm OD}$ and the number of symbols $M_{\rm OD}$ in ODDM are configured as $M_{\rm OD}N_{\rm OD}=N$, so that the two systems transmit $N$ digital symbols with the same subcarrier spacing $\frac{1}{T}$, while occupying nearly the same TF region.
@@
As shown above, AFDM is an $N$-subcarrier system with IFFT complexity $\mathcal{O}(N\log N)$. By comparison, ODDM is an $N_{\rm OD}$-subcarrier system requiring $M_{\rm OD}$ IFFTs of size $N_{\rm OD}$, resulting in a complexity of $M_{\rm OD}\mathcal{O}(N_{\rm OD}\log N_{\rm OD})=\mathcal{O}(N\log N_{\rm OD}),$ which is substantially lower than $\mathcal{O}(N\log N)$, since typically $N_{\rm OD}\ll N$. Furthermore, it is known that a CP is not necessary for ODDM signals [13], as ODDM is essentially a pseudo-impulse-based transmission scheme for DD channels [10]. Consequently, when transmitting the same number of digital symbols with the same subcarrier spacing (frequency resolution), the implementation complexity of ODDM is notably lower than that of AFDM.
@@ VII Conclusion
This paper develops a unified framework for chirp-domain waveforms, including OCDM and AFDM. By examining their continuous-time representations, it is shown that these waveforms naturally fall within the conventional WH framework for multicarrier waveforms, where the root chirp corresponds to the prototype pulse. This perspective further reveals that chirp-domain waveforms can be interpreted as PS-OFDM with a constant-envelope prototype pulse, under which their PSD can be characterized analytically.
Practical implementations based on sub-Nyquist time-domain samples are also analyzed. The resulting aliased waveforms and their orthogonality properties show that orthogonality can be preserved through appropriate sample-wise pulse shaping. Furthermore, the derived I/O relation over DD channels reveals that the effective channel at a practical receiver generally cannot be represented as a superposition of pure path-wise DD components, leading to a non-negligible deviation from the I/O relation commonly used in the literature. Simulations are provided to validate the theoretical results. Future work may extend these findings to receiver design and waveform optimization for off-grid DD channels, as well as to multi-antenna systems.
@@ Appendix A Orthogonality Analysis of Aliased Chirps
$\hat{\phi}_{n}(t)=e^{j2\pi\left(c_{2}n^{2}+\frac{c_{1}}{\Delta t^{2}}t^{2}+\frac{n}{T}t-\frac{q}{\Delta t}t\right)}$ (56)
where $T=N\Delta t$, $C=2N|c_{1}|$, and the interval index $q$ updates at boundaries $t_{n,q}=\frac{T}{C}q-\frac{n}{C}\Delta t$. This boundary condition allows us to express $q$ as an explicit function of $t$, denoted $q_{n}(t)$, using the floor function
@@
[…]
Substitute this into $\Delta q(t)$. Because $\hat{c}$ is an integer, we can use the floor function property $\lfloor x+\hat{c}\rfloor=\lfloor x\rfloor+\hat{c}$ to factor $\hat{c}$ out:
$\Delta q(t)$ $=\left\lfloor(\varepsilon+\hat{c})+\frac{n}{N}\right\rfloor-\left\lfloor(\varepsilon+\hat{c})+\frac{n^{\prime}}{N}\right\rfloor$
[…]
@@
The integer $\hat{c}$ completely cancels out of the expression. Because the resulting expression depends solely on the local time variable $\varepsilon$ and is independent of $\hat{c}$, we introduce the shorthand $\equiv\Delta q(\varepsilon)$. This explicitly denotes that $\Delta q$ is periodic across every segment $\hat{c}$.
Substituting $t=(\varepsilon+\hat{c})\frac{T}{C}$ into the integral and rewriting as a sum over $\hat{c}$, we obtain
$I_{n,n^{\prime}}\propto\frac{T}{C}\sum_{\hat{c}=0}^{C-1}\int_{0}^{1}e^{j2\pi\left(\frac{n^{\prime\prime}}{T}-\frac{\Delta q(\varepsilon)}{\Delta t}\right)(\varepsilon+\hat{c})\frac{T}{C}}d\varepsilon.$ (60)
[…]
$I_{n,n^{\prime}}\propto\frac{T}{C}\int_{0}^{1}e^{j2\pi\frac{\eta(\varepsilon)}{C}\varepsilon}\underbrace{\left[\sum_{\hat{c}=0}^{C-1}e^{j2\pi\frac{\eta(\varepsilon)}{C}\hat{c}}\right]}_{\text{Geometric Sum }S(\varepsilon)}d\varepsilon,$ (61)
@@
Let us define an $N\times N$ matrix $\mathbf{I}$ with entries $I_{n,n^{\prime}}$. The simulation results for $|I_{n,n^{\prime}}|$ under $N=32$ are provided below to verify the conditional orthogonality of the aliased chirps. Figs. 5 and 6 show the results for $C=48$ and $C=32$, respectively, corresponding to Case 1 ($C\geq N$). As observed, the matrix $\mathbf{I}$ is almost entirely concentrated on the main diagonal, indicating that a significant inner product is obtained only when $n=n^{\prime}$. This confirms that distinct aliased chirps remain mutually orthogonal, as predicted for Case 1. In contrast, Fig. 7 displays the results for $C=16$, corresponding to Case 2 ($C<N$). Here, $\mathbf{I}$ exhibits two distinct off-diagonal “bright bands” at $|n-n^{\prime}|=16$, exactly matching the chirp-index separation specified by $C$. This loss of orthogonality occurs at a structured offset rather than being randomly distributed, aligning with the theoretical predictions for Case 2.
Figure 5: $|I_{n,n^{\prime}}|$, $C=48$, $N=32$.
Figure 6: $|I_{n,n^{\prime}}|$, $C=32$, $N=32$.
Figure 7: $|I_{n,n^{\prime}}|$, $C=16$, $N=32$.